Problem: There are three consecutive even integers. Half the sum of the second and third numbers is 35. What are the integers? I know the answer is 32, 34, 36, but I don't know how to write the equation. Your site has helped me solve other problems. Thank you.
Answer by stanbon(75887)

There are three consecutive even integers. Half the sum of the second and third numbers is 35. What are the integers?
---------------------------
"Even" means a multiple of two.
---------------------------
1st: 2x
2nd: 2x+2
3rd: 2x+4
-----------
EQUATION:
(1/2)(2x+2+2x+4)= 35
4x+6 = 70
4x = 64
x = 16
---------------
2x = 32
2x+2 =34
2x+4 = 36
=====================
Cheers,
Stan H.